OpenStudy (anonymous):

Line through (-3, -4) and slope 2/3 This equation has to be written in Slope-Intercept form. CAN ANYONE HELP?

OpenStudy (mathmale):

You are given a point through which the line passes and its slope. Insert this info into the point-slope form of the equation of a straight line. Once you've done that, simplify the result to obtain the equation of the line in slope-intercept form.

OpenStudy (whpalmer4):

point-slope form for a line with slope \(m\) passing through point \((x_0,y_0)\) is \[y-y_0 = m(x-x_0)\] slope-intercept form is \[y = mx+b\] You can get slope-intercept form from point-slope form by solving for \(y\)